What mass of sodium hydroxide can be made by reacting 2.3 grams of sodium with water


1
Expert's answer
2021-03-02T01:59:32-0500

2Na + 2H2O = 2NaOH + H2

n(NaOH) = n(Na) = (2.3 g)/(23 g/mol) = 0.1 mol

m(NaOH) = (0.1 mol)*(40 g/mol) = 4 grams